Find a matrix A whose minimal polynomial is t^3-5t^2+6t+8
1
Expert's answer
2020-12-21T17:59:26-0500

The matrix associated to a generic minimal polynomial "p(x)=a_0+a_1x+a_2x^2+\cdots +a_{n-1}x^{n-1}+a_nx^n" is given as;


"A=\begin{pmatrix}
 0&-a_0 \\
 I&-a
\end{pmatrix}"

Where "I" is the "(n-1) \times (n-1)" identity matrix and "a=(a_1,\cdots, a_{n-1})^T". A is an "n\times n" matrix.


So, for the question,

"a_0=8,a=(6,-5)^T, I=\begin{pmatrix}
 1&0\\
 0&1
\end{pmatrix}"


"A=\begin{pmatrix}
 0&0&-8\\
 1&0&-6\\
 0&1&5\\
\end{pmatrix}"